OpenStudy (anonymous):

Rewrite the expression in terms of the given function or functions. (sec x + csc x) (sin x + cos x) - 2 - cot x; tan x A. 0 B. 2tan x C. tan x D. 2 + tan x

OpenStudy (anonymous):

I got C is that right?

OpenStudy (helder_edwin):

is the expression like this \[ \large (\sec x+\csc x)(\sin x+\cos x)-2-\cot x/\tan x \]

OpenStudy (campbell_st):

well rewrite the brackets and simply that 1st \[(\frac{1}{\cos(x)} + \frac{1}{\sin(x)})\times (\sin(x) + \cos(x)) \] \[\frac{\sin(x)}{\cos(x)} + 1 + 1 + \frac{\cos(x)}{\sin(x)}\] simplifies to \[\tan(x) + \cot(x) + 2\] so the equation becomes \[\tan(x) + \cot(x) + 2 - 2 -\cot(x) = \tan(x)\] hope it helps

OpenStudy (anonymous):

thnx!

OpenStudy (anonymous):

wait i dont understand the second step..

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

the expansion...?

OpenStudy (anonymous):

sin(x)/cos(x)+1+1+cos(x)/sin(x)

OpenStudy (anonymous):

@campbell_st

OpenStudy (anonymous):

can someone explain the step to me?

OpenStudy (anonymous):

just expand the previous line... it's like FOILing....

OpenStudy (anonymous):

ohh i get it! thnx @dpaInc

OpenStudy (campbell_st):

its just the multiplication 1/sin(x) * sin(x) = 1 1/cos(x) * cos(x) = 1 1/cos(x) * sin(x) = sin(x)/cos(x) = tan(x) 1/sin(x) * cos(x) = cos(x)/sin(x) = cot(x)

OpenStudy (anonymous):

yw...:)